It has been quiet during the day at the HET this week with day staff on holiday. If the weather clears the night staff will be plenty busy working through the queue. The current weather outlook for the week is cloudy with possible showers. Monday night we were able to do science through thin clouds and successfully collected data for many programs that do not require perfect seeing or zero cloud cover. The bulk of the science data collected in the beginning of the week has been with LRS2-B and LRS2-R. Fingers crossed the weather clears as we approach the new moon on Thursday.
This week marks the first full week of the new trimester (December – March). During this period we will be observing for 2 weeks around the new moon with LRS2-B and LRS2-R and to a very limited extent the first 16 units of VIRUS for science projects from our partners. We have 27 separate projects from our 4 partners. During this first week of the new trimester we took data for 10 programs.
During the 2 weeks around the full moon we will be concentrating on engineering but we may if it does not impose any constraint on the engineering obtain a few science targets during grey and bright time.
It is good to be back online taking science data.
This week has been a busy one for the virus units at the HET. All of the virus units were removed from the enclosure, backfilled, pumped to vacuum and reinstalled. While doing this cooling fans were installed on all of the virus controllers to help keep the electronics cool. They hope that this will make the virus system a bit more stable.
The HET has been churning out lots of science during dark time this month as well. All groups that have submitted targets have received data!
In the previous week the HET Board of Directors met and were generally happy with the progress being made on the HET. They are very eager for science operations to begin and have authorized the start of Early Science Operations starting July 1st. This will mean that HET member institutions will be able to submit requests to their local Telescope Allocation Committee in the coming weeks. The amount of time available will grow with time but is likely to start with 1 week per dark period. The instrument that will be available for Early Science Operations will be the new Low Resolution Spectrograph (LRS2).
In this week the HET continued with VIRUS commissioning during bright time with the addition of 7 new VIRUS units. This brings the number of spectrographs units (and IFUs) to 16. We will use these during commissioning in the coming dark run but we can already see that several of them will need some further engineering while others are very good.
The mirror team continues to swap mirrors into the array this week they swapped out 4 mirrors. The array now has an average coating age of 454 days with a swap rate of 1.7 segments/week.
We are currently in bright time (near full moon) and the currently active instruments, LRS2 and VIRUS are really designed to operate in dark time (when the moon is below the horizon), thus we spent this week analyzing the data taken in the last dark run, testing and updating the telescope control software and training the night time staff to make sure all of the staff have the same skills and knowledge of the system for the upcoming dark run.
One of the problems we were having in the last dark run was a mysterious end of track due to a hexapod hitting the software limit. Further investigation has revealed that this limit is part of the natural extension of the warming tracker. As the steel structure expands from warmer temperatures we have to move the hexapods down to compensate. We also found that we had set the software limits fairly far from the limit switches so we could give ourselves plenty more room without any changes to the hardware of the telescope. This is an example of the type of tuning that we continue to have to do with the control algorithms.
Both the LRS and the VIRUS are IFU (Integral Field Unit) spectrographs, which means that one could take the spectra for each fiber and reconstruct an image of the field which was observed. Below is the very first reconstruction that we did for a VIRUS field. It was a fairly bright (8th mag) standard star. The image scale is set very wide so that we can see all of the interesting features that come from the unusual design of the HET. The six image spikes comes from the edges of the hexagonal HET segments, and the stronger spikes along the X direction come from the tracker Y superstructure.
Not long after generating this the software team was able to reconstruct a number of images for each IFU in a pointing of a globular cluster, NGC5272. Below I show a couple of examples of these.
Here is a schematic of what the current array of IFUs would look like compared to a Digitized Sky Survey Image of the field we pointed at. This is a by-product of our current setup software.
We have had a very exciting couple of weeks. In the last blog post I had mentioned that we made some repairs and improvements to the top end of the telescope and that the LRS2 team had left the telescope in the hands of the night operations team to gain experience with the system. That is exactly what we did over the bright lunation. We had nights set aside with all of the telescope operators working together at night to make sure that we all know how to operate all of the new software the same way and to work out the kinks in the operations procedures. This was followed by a few nights with the resident astronomers all working at night to do the same. The result was that we had developed a few pieces of code that would allow us to determine the exact position of the LRS2 field of view and could setup any target that the Acquisition camera could see.
In the mean time, the day staff was prepping the right VIRUS enclosure for the delivery of the first 9 units (18 spectrographs). The spectrographs were brought out in two vans along with the VIRUS commissioning team for the dark run (where the moon is in the new phase). The fibers were strung, the units were vacuum pumped and finally cooled. Everything went ahead of schedule and on Saturday night we were able to make use of some very clear skies to get first light with the VIRUS spectrographs. On Sunday night we were able to get the first LRS2 science spectrum (a supernova for an astronomer at PSU) and a few long exposures for characterization of the VIRUS units.
The telescope is still occasionally fussy but generally performed well and the commissioning work continues on into the next week until quarter moon. At the present time the data reduction software is the thing that is lagging most behind.
Two weeks ago was the HET Board of Director’s meeting in Penn State. The meeting lasted two days and the Board got status reports from HET operations and each of the instrument teams. The main news is that that LRS2 is expecting to reach first science in the first few months of 2016, HRS2 is going to start commissioning before Summer 2016 and VIRUS units are going to be coming in over the next 9 months. The Board was encouraged by the progress being made and hopes that we can continue the pace. They were also pleased to see the progress being made on HPF and were impressed with the clean room facility tour they were given. No major changes or action items were reported by the Board.
Once the commissioning team returned from Happy Valley, we went right back to work and were able to push along the closure of one of our major metrology loops, the guide probes. I am pleased to report that we are able to guide at any telescope Az for full trajectories with the probes at any position within their range.
This week we had lots of visitors from Austin. We had new equipment to install on the tracker, for example, the field calibration unit. This is the device at the entrance aperture of the corrector which can shine light from flat lamps or from line lamps for calibration of our spectrographs.
We also had delivery of the LRS2-R spectrograph and a temporary installation of a VIRUS unit. The VIRUS unit was put in place mostly to test the mechanics of the enclosure cooling and control systems. We were able to get our first spectrum with both spectrographs. By the end of the week we were able to find a guide star with guide probes at one specific orientation and then hold a science target on a specific location of the LRS2-R and VIRUS IFUs. We then could move the target around the IFU to confirm the direction within the IFU. Even the weather cooperated finally, with some near photometric conditions we put a spectrophotometric standard on the LRS2-R. On the VIRUS unit we confirmed that the dither mechanism does work and we were able to move a target from one fiber to the adjacent fiber.
We will be moving back to guide probe commissioning in the coming week and may not get back to LRS2 instrument commissioning until January.
All in all a very exhausting, but good week!!!
We did some simulations of data transfer rates in preparation for VIRUS coming on line. The data set consists of 900 fits files; each file is 8524800 bytes in size. The working assumption is that one Hetdex image consists of 150 spectrographs x 2 amplifiers/spectrograph x 3 dither per field with a CCD size of 1032×2064 pixels. Total data size is 7,672,320,000 bytes (7.3 GB). The data would be collected over 20 minutes (~6 minutes per dither) but it only took 449.5 seconds (7.5 minutes) to transfer this data to Austin. Thus we should be able to keep up with the data collection rate without bringing the network to a crawl.
The large air and glycol (“Blue Thunder”) hoses were installed through the Az wrap and laid in the trench in the floor and through one of the ports through the pier to the ring wall in the last two weeks. The next step will be to connect these to the outside lines and do a pressure test of the system.
All power connections for the VIRUS enclosures are now complete and tested. We have also conducted some heat load tests on the liquid nitrogen system by placing plastic bags over each of the VIRUS bayonets (the component that will stick into a VIRUS unit to keep it cold). So far most of the system checks out close to specification.
The primary mirror group installed 4 mirrors this week and continue to keep up their 30% over expected swap rate in hopes of being close to completion when the corrector is ready for commissioning.
The CCAS shutter failed this week and could not be repaired from inside the CCAS Dome. Instead of having to rent a very expensive 90 ft manlift the mechanical team found a way to safely undo the mounting to the CCAS Dome and lower it to the ground. A replacement shutter is on order and should arrive in the next week or two. It will be installed by just reversing the procedure developed to remove the old shutter. Until then we’ve installed a simple wood plug that can be removed manually (after climbing the 90ft tower) to continue with critical primary mirror work.
Happy New Year! The blogger is back from Winter break and travels to January Science meetings.
We will start with the bad news. The tests on the corrector have determined that there is significant asymmetries. These are asymmetries beyond what were found and reported to the board of directors in the December meeting. Those aberrations were going to be removed by changing the plate glass that was going in to seal the bottom of the corrector into a fifth optical element with some small power. That optic has been fabricated and installed in the corrector. The newly discovered astigmatism was found off axis and might have an impact on image quality at the edge of the field of view. To further investigate this optical issue we are adding a few months to the delivery date of the corrector. Instead of being delivered in late January the new delivery date is expected to be early May and significant on-sky commissioning during the rainy season perhaps pushing us into September.
On a more positive note the 2nd VIRUS enclosure has arrived and installed.
There is a lot of plumbing and electrical work required to finish the installation. This is our highest priority.
During the long break the Remote Thermal Area contractors have installed the glycol chiller and are in the process of commissioning it with glycol. This nearly completes the Remote Thermal Area project. Only a few punch list items remain.
The HPF doors also arrived and have been installed. The HPF is now thermally isolated from the spectrograph room. The Penn State Team will monitor the temperatures inside and outside of their enclosure to see if our new Mitsubishi units can hold the temperature to a tight enough tolerance for their specifications. Meanwhile the insulating panels for the HRS II have arrived and will be installed in the coming month. These panels along with an active feed back heater inside the enclosure should allow the HRS II to be held to extremely tight tolerances, < 0.1 C.
The mirror team has started mirror swaps and started using the Strip and Wash room to remove old mirror coatings and prep them for the coating chamber. Our goal is to get the team up to 4 mirrors swaps every 3 weeks.